OpenStudy (anonymous):

Solve x2 = 144. 72 12 ±12 −12

OpenStudy (freckles):

well what numbers squared give you 144?

OpenStudy (freckles):

here is example: solve x^2=9 well 3^2 and (-3)^2 both give by 9 so x=3 or x=-3 are both solutions to x^2=9

OpenStudy (freckles):

there are two solutions if you take the square of both sides you need to be careful of how you interpret sqrt(x^2) this actually is the same thing as |x| so \[x^2=144 \\ \sqrt{x^2}=\sqrt{144} \\ |x|=12 \\ \] this will give two solutions (not just one)

OpenStudy (anonymous):

Thanks

OpenStudy (condor34):

x= positive or negative 12